Constants Previous Answers An electrically neutral thin, square metal slab measuring 4.00 on a side, is placed in a uniform external field that is perpendicular to its square area. VCorrect ▼ Part C If the external field strength is 1280 N/C, what is the strength of the field that is generated by the charges induced on the slab? Eg= 2.05 N/C Submit Previous Answers Request Answer XIncorrect; Try Again; 4 attempts remaining
Answer #1

Solution:

Electric field strength due to the slab will be Zero.

As we know that, electric stength of a slab with uniform surface charge density \sigma is given by

E=\frac{\sigma }{\varepsilon _{o}}

But, in the current case, the slab is neutral and charges cann't be induced with the external electric field.

So,

\sigma =0

Therefore,

E=\frac{\sigma }{\varepsilon _{o}}=\frac{0}{\varepsilon _{o}}=0